# Taming Randomness in Agent-Based Models using Common Random Numbers

### v1

@@ Abstract
Random numbers are at the heart of every agent-based model (ABM) of health and disease. By representing each individual in a synthetic population, agent-based models enable detailed analysis of intervention impact and parameter sensitivity. Yet agent-based modeling has a fundamental signal-to-noise problem, in which small differences between simulations cannot be reliably differentiated from stochastic noise resulting from misaligned random number realizations. We introduce a novel methodology that eliminates noise due to misaligned random numbers, a first for agent-based modeling. Our approach enables meaningful individual-level analysis between ABM scenarios because all differences are driven by mechanistic effects rather than random number noise. A key result is that many fewer simulations are needed for some applications. We demonstrate the benefits of our approach on three disparate examples and discuss limitations.
keywords:
Agent-based Model, Simulation, Random Numbers, Epidemiology
1 Introduction
@@ 2 Methods
2.1 Separate pseudo-random number streams for each decision
@@
2.2 Time step-dependent PRNG stream jumping
@@
2.3 At most one call to each PRNG per time step
@@
2.4 Slot-based assignment of random number draws
@@
2.5 Pairwise random numbers
@@
After exploring several alternatives and checking for bias, see Appendix A, we settled on the following transformation,
@@
2.6 Network edge formation
@@
2.6.1 Erdős-Rényi Dynamic Random Network
In an Erdős-Rényi graph, each pair of nodes is connected with probability $p$. Ordinarily, $N(N-1)/2$ random numbers would be used in assessing the existence of an undirected edge between each pair of agents. To create a CRN-safe Erdős-Rényi network, we instead use one of the pairwise random number methods described in Section 2.5 (other than Modulo, which is biased). With this approach, edges can have a defined duration or be can recreated on each time step. The loss or addition of agents will not affect other network edges. We note that this approach could be generalized. For example, the probability, $p$, of an edge could depend on agent properties, simulation time, or other factors.
2.6.2 Dynamic Disk Graph
@@ 2.7 Pathogen transmission
We have identified several solutions to overcome this challenge. We describe an acquisition-based approach in Appendix B. The primary approach we have implemented takes advantage of the pair-specific random numbers, as described in Section 2.5. This method works just like the naïve approach, but substitutes a pair-specific random realization instead of a random draw from a centralized generator.
3 Results
@@
All results come from Starsim [14] (v1.0), an open-source general-purpose health and disease modeling framework. The methods to achieve CRN, as described in Section 2, have been designed into the Starsim framework. Simultaneously, the framework features the ability to disable CRN, reverting back to a single centralized random number generator for comparison purposes. We present results comparing the following two approaches to generating random numbers.
@@
3.1 Maternal & Child Health: Prevention of postpartum hemorrhage
@@
To explore the potential benefits of averting PPH, we apply our approach to common random numbers on a synthetic population resembling sub-Saharan Africa. These results demonstrate the ability of the CRN-based approach to simulate vital dynamics, births and deaths, as made possible by the “slots” described in Section 2.4.
@@
Figure 1: Results from PPH simulations comparing centralized (left) and common random number (right) approaches. The top row shows the absolute number of maternal deaths occurring in this synthetic population with the shaded regions corresponding to plus or minus one standard deviation. The second row shows differences between each of the two coverage levels and the reference scenario, paired by random number seed. The third row shows the time-evolution of the Pearson correlation coefficient for three output channels, as indicated by linetype. Finally, the bottom row shows the distribution of differences in maternal deaths at the the final time. Colors indicate coverage of the PPH prevention intervention ranging from 0% (Reference, green) to 10% (red) and 90% (blue).
@@
Additional correlation results for this PPH example can be found in Appendix D.
3.2 Infectious disease transmission
@@
Simulation results presented in this section use SIR dynamics with an exponentially-distributed duration of infection with a mean of 30 days. The infection fatality ratio is set at 5%. We default to a Barabási–Albert “power law” network topology with parameter $m=1$. Static networks are CRN-safe, recall Section 2.6.4. We select a time step of one day and set the default transmissibility parameter, $\beta$, so that there is a 20% chance of transmission per day between each connected pair of infected an susceptible agents. The size of the simulated population is varied. Simulations start on the first day of 2020 and end 6-months later. The simulated vaccine, which acts to reduce susceptibility to infection acquisition, is distributed on day five at 5% or 90% coverage and is assumed to have a constant efficacy of 70%.
@@
Figure 2: Timeseries result for the SIR example for population sizes ranging from 10 (red) to 1,000 (purple) for centralized (left) and CRN (right) approaches to random number generation. The top row shows the cumulative number of infections as a percentage of the population whereas the middle row shows the difference in cumulative incidence between the 90% coverage level and the no-vaccine reference scenario. The final row shows how the Pearson correlation coefficient against the reference evolves over time for cumulative incidence with 5% (solid) and 90% (dashed) vaccine coverage.
@@
Additional results generated by varying the topology of the SIR network are presented in Appendix E.
3.3 HIV prevention via voluntary medical male circumcision
@@
Figure 3: Results from the voluntary medical male circumcision example. The top row shows HIV prevalence (left) and VMMC coverage in men (right) in the 15-49 age group. Link color distinguishes scenario between continuing the VMMC program (red) or discontinuing in 2020 (blue). The middle panel shows the fold reduction in the number of simulation replicates that would be required to achieve a given standard error at each point in time during the simulation. Infections averted and deaths averted are displayed in blue and purple, respectively. Dashed vertical lines mark 5, 20, and 50 year horizons. Finally, the bottom row shows cumulative distributions of infections averted 5-, 20-, and 50-years after the beginning of the intervention. Results are presented for the centralized (green) and CRN (pink) generators.
@@ 4 Discussion
Our approach has several limitations. First, model re-engineering may be required to retrofit an existing agent-based model with the methods described in Section 2. Second, the resulting model code could be more challenging to use and modify. In traditional modeling, a user could call the system “rand” function, which accesses a centralized generator; here additional care must be taken by users when evaluating stochastic decisions.
While random number generation typically takes a small percentage of overall simulation time, our slot-based approach to random numbers draws many more random numbers than are actually used. While this seems wasteful and does reduce model performance, random number generation is not a significant performance bottleneck in our experience. Instead, we have found the embedding network has $O(N^{3})$ scaling that dramatically affects performance with large $N$. Users seeking more performance or larger population sizes could use the disk or Erdős-Rényi algorithms we described. The performance of these algorithms is assessed in Appendix C.
@@
[…]
Software and analysis code availability Methods described in this article have been implemented in Starsim , which is available as open source code on GitHub [14]. Results were generated using v1.0.1 of the framework. Analysis code is available GitHub repository available online at https://github.com/starsimhub/crn_paper.
[…]
Acknowledgments
Appendix A Pairwise random numbers
@@
Figure 4: Frequencies of random graphs on $N=4$ nodes created using each pairwise approach to random number generation against the “True Random” pseudo-random reference. Each hash code on the X-axis corresponds to one unique possible graph. Bar heights show the number of times each graph occurred in two-million replicates. Bar color indexes the method of random number generation used in creating edges.
@@
Appendix B Acquisition-based disease transmission
@@
Appendix C Network performance characterization
@@
Figure 5: Performance scaling of CRN-safe network generation algorithms. We plot the time to generate a network, measured in seconds, as a function of the number of agents from 10 to 32,000. Line color indicates the network generation algorithm. Of the algorithms compared, the Random network is the only one that is not CRN safe. Note the log10 scale on both X- and Y-axes.
Appendix D Additional PPH correlation results
@@
Figure 6: Scatter plot comparing cumulative counts of births (top row) and maternal deaths (bottom row) for 10% (left) and 90% (right) coverage levels. Individual points represent seed-matched pairs of simulations, the reference result on X and the intervention result on Y. Dot color and shape reference the random generator used.
Appendix E Varying the SIR network topology
@@
Figure 7: Pearson correlation of the number of recovered agents between reference and vaccine scenarios for centralized (left) and CRN (right) approaches. Each line color represents one network topology, as indicated in the legend, and the line style indicates coverage levels of 5% (solid) and 80% (dashed).

### v2

# Noise-free comparison of stochastic agent-based simulations using common random numbers

@@ Abstract
Random numbers are at the heart of every agent-based model (ABM) of health and disease. By representing each individual in a synthetic population, agent-based models enable detailed analysis of intervention impact and parameter sensitivity. Yet agent-based modeling has a fundamental signal-to-noise problem, in which small changes between simulations cannot be reliably differentiated from stochastic noise resulting from misaligned random number realizations. We introduce a novel methodology that eliminates noise due to misaligned random numbers, a first for agent-based modeling. Our approach enables meaningful individual-level analysis between ABM scenarios because all differences are driven by mechanistic effects rather than random number noise. We demonstrate the benefits of our approach on three disparate examples. Results consistently show reductions in the number of simulations required to achieve a given standard error with levels exceeding 10-fold for some applications.
Author summary
We present new computational methodology that addresses a longstanding signal-to-noise problem in agent-based modeling that arises when comparing simulation outcomes. With the traditional approach that we and other modelers have used for decades, random draw misalignment between simulations results in high variance and implausible differences, complicating impact evaluation, parametric sensitivity, and scenario analysis. Our new method achieve perfect alignment of random draws between simulations, thereby preventing stochastic branching entirely. Similar ideas have been demonstrated for simple cohort models, but those techniques did not work for key aspects we need in disease modeling like dynamic populations, births and in-migration specifically, and agent-to-agent interactions, as needed for pathogen transmission. We tested our new methodology on three use cases and found it has many benefits including dramatic reductions in the number of simulation replicates required for some applications. We believe that practitioners both within and beyond the field of computational epidemiology will benefit considerably from this improved approach to agent-based modeling.
Introduction
@@
Separate pseudo-random number streams for each decision
@@
Time step-dependent PRNG stream jumping
@@
At most one call to each PRNG per time step
@@
Slot-based assignment of random number draws
@@
Pairwise random numbers
@@
After exploring several alternatives and checking for bias, see S1 Appendix, we settled on the following transformation,
@@
Network edge formation
@@
Erdős-Rényi Dynamic Random Network
In an Erdős-Rényi graph, each pair of nodes is connected with probability $p$. Ordinarily, $N(N-1)/2$ random numbers would be used in assessing the existence of an undirected edge between each pair of agents. To create a CRN-safe Erdős-Rényi network, we instead use one of the pairwise random number methods described in Pairwise random numbers (other than Modulo, which is biased). With this approach, edges can have a defined duration or be can recreated on each time step. The loss or addition of agents will not affect other network edges. We note that this approach could be generalized. For example, the probability, $p$, of an edge could depend on agent properties, simulation time, or other factors.
Dynamic Disk Graph
@@
We have identified several solutions to overcome this challenge. We describe an acquisition-based approach in S2 Appendix. The primary approach we have implemented takes advantage of the pair-specific random numbers, as described in Pairwise random numbers. This method works just like the naïve approach, but substitutes a pair-specific random realization instead of a random draw from a centralized generator.
Results
@@
All results come from Starsim [14] (v1.0), an open-source general-purpose health and disease modeling framework. The methods to achieve CRN, as described in Materials and methods, have been designed into the Starsim framework. Simultaneously, the framework features the ability to disable CRN, reverting back to a single centralized random number generator for comparison purposes. We present results comparing the following two approaches to generating random numbers.
@@
Maternal & Child Health: Prevention of postpartum hemorrhage
@@
To explore the potential benefits of averting PPH, we apply our approach to common random numbers on a synthetic population resembling sub-Saharan Africa. These results demonstrate the ability of the CRN-based approach to simulate vital dynamics, births and deaths, as made possible by the “slots” described in Slot-based assignment of random number draws.
@@
Fig 1: Results from PPH simulations comparing centralized (left) and common random number (right) approaches. The top row shows the absolute number of maternal deaths occurring in this synthetic population with the shaded regions corresponding to plus or minus one standard deviation. The second row shows differences between each of the two coverage levels and the reference scenario, paired by random number seed. The third row shows the time-evolution of the Pearson correlation coefficient for three output channels, as indicated by linetype. Finally, the bottom row shows the distribution of differences in maternal deaths at the the final time. Colors indicate coverage of the PPH prevention intervention ranging from 0% (Reference, green) to 10% (red) and 90% (blue).
@@
Additional correlation results for this PPH example can be found in S4 Appendix.
Infectious disease transmission
@@
Simulation results presented in this section use SIR dynamics with an exponentially-distributed duration of infection with a mean of 30 days. The infection fatality ratio is set at 5%. We default to a Barabási–Albert “power law” network topology with parameter $m=1$. Static networks are CRN-safe, recall Static network. We select a time step of one day and set the default transmissibility parameter, $\beta$, so that there is a 20% chance of transmission per day between each connected pair of infected an susceptible agents. The size of the simulated population is varied. Simulations start on the first day of 2020 and end 6-months later. The simulated vaccine, which acts to reduce susceptibility to infection acquisition, is distributed on day five at 5% or 90% coverage and is assumed to have a constant efficacy of 70%.
@@
Fig 2: Timeseries result for the SIR example for population sizes ranging from 10 (red) to 1,000 (purple) for centralized (left) and CRN (right) approaches to random number generation. The top row shows the cumulative number of infections as a percentage of the population whereas the middle row shows the difference in cumulative incidence between the 90% coverage level and the no-vaccine reference scenario. The final row shows how the Pearson correlation coefficient against the reference evolves over time for cumulative incidence with 5% (solid) and 90% (dashed) vaccine coverage.
@@
Additional results generated by varying the topology of the SIR network are presented in S5 Appendix.
HIV prevention via voluntary medical male circumcision
@@
Fig 3: Results from the voluntary medical male circumcision example. The top row shows HIV prevalence (left) and VMMC coverage in men (right) in the 15-49 age group. Link color distinguishes scenario between continuing the VMMC program (red) or discontinuing in 2020 (blue). The middle panel shows the fold reduction in the number of simulation replicates that would be required to achieve a given standard error at each point in time during the simulation. Infections averted and deaths averted are displayed in blue and purple, respectively. Dashed vertical lines mark 5, 20, and 50 year horizons. Finally, the bottom row shows cumulative distributions of infections averted 5-, 20-, and 50-years after the beginning of the intervention. Results are presented for the centralized (green) and CRN (pink) generators.
@@ Discussion
Our approach has several limitations. First, model re-engineering may be required to retrofit an existing agent-based model with the methods described in Materials and methods. Second, the resulting model code could be more challenging to use and modify. In traditional modeling, a user could call the system “rand” function, which accesses a centralized generator; here additional care must be taken by users when evaluating stochastic decisions.
While random number generation typically takes a small percentage of overall simulation time, our slot-based approach to random numbers draws many more random numbers than are actually used. While this seems wasteful and does reduce model performance, random number generation is not a significant performance bottleneck in our experience. Instead, we have found the embedding network has $O(N^{3})$ scaling that dramatically affects performance with large $N$. Users seeking more performance or larger population sizes could use the disk or Erdős-Rényi algorithms we described. The performance of these algorithms is assessed in S3 Appendix.
@@
Supporting information
Software and analysis code availability
Methods described in this article have been implemented in Starsim , which is available as open source code on GitHub [14]. Results were generated using v1.0.1 of the framework. Analysis code is available GitHub repository available online at https://github.com/starsimhub/crn_paper.
S1 Appendix
Pairwise random numbers
@@
Fig 4: Frequencies of random graphs on $N=4$ nodes created using each pairwise approach to random number generation against the “True Random” pseudo-random reference. Each hash code on the X-axis corresponds to one unique possible graph. Bar heights show the number of times each graph occurred in two-million replicates. Bar color indexes the method of random number generation used in creating edges.
@@
S2 Appendix
Acquisition-based disease transmission
@@
S3 Appendix
Network performance characterization
@@
Fig 5: Performance scaling of CRN-safe network generation algorithms. We plot the time to generate a network, measured in seconds, as a function of the number of agents from 10 to 32,000. Line color indicates the network generation algorithm. Of the algorithms compared, the Random network is the only one that is not CRN safe. Note the log10 scale on both X- and Y-axes.
@@
S4 Appendix
Additional PPH correlation results
@@
Fig 6: Scatter plot comparing cumulative counts of births (top row) and maternal deaths (bottom row) for 10% (left) and 90% (right) coverage levels. Individual points represent seed-matched pairs of simulations, the reference result on X and the intervention result on Y. Dot color and shape reference the random generator used.
@@
S5 Appendix
Varying the SIR network topology
@@
Fig 7: Pearson correlation of the number of recovered agents between reference and vaccine scenarios for centralized (left) and CRN (right) approaches. Each line color represents one network topology, as indicated in the legend, and the line style indicates coverage levels of 5% (solid) and 80% (dashed).